OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x=4y 3x-2y=20

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

helppppppppppppppppp

OpenStudy (anonymous):

3(4y) - 2y = 20 y = 2 so x = 4*2 = 8 answer (8,2)